[Hallc_running] [New Logentry] RC Report - 8/26/22

arunts at jlab.org arunts at jlab.org
Fri Aug 26 11:00:02 EDT 2022

Logentry Text:
We changed to negative polarity runs during day shift yesterday and started with 7 uA beam current. We scanned a couple of beam current settings for ~20 mins (7, 10, 15, 30, 25 uA) and settled on 11 uA for the day shift. Eventually after looking at the tracking efficiencies, we settled at 12 uA for the rest of the swing and owl shifts (https://logbooks.jlab.org/entry/4033932). The plan is to continue data taking at this configuration for another ~2.5 days. For the small angle configuration change that happens sometime during the weekend, Steve Lassiter is happy to monitor this configuration change remotely so we DO NOT need an access. 

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071

Important studies:
Rates @ 12uA: LD- Q2=3.85 Thshms=8.000deg: https://logbooks.jlab.org/entry/4033868
Rates @ 11uA: LD- Q2=3.85 Thshms=8.000deg: https://logbooks.jlab.org/entry/4033905
Examining event cuts, tracking and detector efficiency calculations: https://logbooks.jlab.org/entry/4033995
Quick Check Pi- Per Unit Charge at LD- Q2=3.85 THshms=8.00deg setting: https://logbooks.jlab.org/entry/4033959


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4034120
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220826/5401007c/attachment.html>

More information about the Hallc_running mailing list